# Corresponding author(s)

### v1

@@ Abstract
Recent advancements in slow-thinking reasoning models have shown exceptional performance in complex reasoning tasks. However, these models often exhibit overthinking-generating redundant reasoning steps for simple problems, leading to excessive computational resource usage. While current mitigation strategies uniformly reduce reasoning tokens, they risk degrading performance on challenging tasks that require extended reasoning. This paper introduces Difficulty-Adaptive Slow-Thinking (DAST), a novel framework that enables models to autonomously adjust the length of Chain-of-Thought(CoT) based on problem difficulty. We first propose a Token Length Budget (TLB) metric to quantify difficulty, then leverage length-aware reward shaping and length preference optimization to implement DAST. DAST penalizes overlong responses for simple tasks while incentivizing sufficient reasoning for complex problems. Experiments on diverse datasets and model scales demonstrate that DAST effectively mitigates overthinking (reducing token usage by over 30% on average) while preserving reasoning accuracy on complex problems.
@@ 1 Introduction
Recently, significant advancements have been made in slow-thinking reasoning models, exemplified by OpenAI’s o1[11] and DeepSeek-R1[2]. These slow-thinking reasoning models, which simulate human deep-thinking mechanisms through self-reflection, error correction, and exploration of multiple solution strategies, have demonstrated remarkable potential in complex reasoning tasks such as mathematical problem-solving[8] and programming challenges[4].
Figure 1: An Example of overthinking phenomenon from DeepSeek-R1.
@@
However, empirical studies [1, 15] have shown that these reasoning models suffer from overthinking phenomenon. In other words, these models tend to generate redundant solutions and unnecessarily complex reasoning steps when addressing simple problems, leading to inefficient computational resource utilization. For instance, as demonstrated in Figure 1, traditional LLM can solve basic mathematic problems such as “3x + 7=22, x=?” with just 58 tokens, while reasoning model such as DeepSeek-R1[2] may consume over 1000 tokens for the same problem. This overthinking phenomenon not only significantly reduces reasoning efficiency but also causes information overload for the users.
@@
Current approaches [1, 14] to mitigate the overthinking problem typically employ a one-size-fits-all strategy, uniformly reducing reasoning steps or token counts across all problems. Although these approaches significantly reduce the output length of slow-thinking models, they risk causing performance degradation, particularly when addressing challenging problems. Prior studies [15, 10] have demonstrated that adequate reasoning length is critical for slow-thinking models to effectively solve complex tasks. It is therefore essential to devise approaches that mitigate overthinking phenomena while maximally preserving reasoning capabilities.
@@
This raises a fundamental question: Can we develop solutions that enable slow-thinking models to autonomously determine reasoning depth according to problem difficulty, thereby generating concise responses for simple questions while maintaining sufficiently extended chain-of-thought reasoning for complex problems? We propose a Difficulty-Adaptive Slow-Thinking (DAST) framework to address this challenge.
@@
Our key idea is straightforward: given that tasks of varying difficulty levels inherently demand different reasoning depths, we propose to establish a mapping relationship between problem complexity and target response length. By comparing the length of the current response with the target response length, we can determine whether to apply additional rewards or penalties to the current answer. Building upon this, we construct a training objective to achieve adaptive reasoning. Specifically, we first introduce a difficulty quantification metric termed ”Token Length Budget”(TLB), which integrates both the accuracy of sampled responses and their length distributions. This metric effectively combines problem difficulty characteristics with token length information. For multiple generated responses sampled, our method applies length-aware reward shaping: Responses exceeding the TLB of simple questions receive penalty signals, while those approaching the TLB for complex problems receive positive incentives. This mechanism allows us to construct pair-wise length preference training data that inherently encodes the relationship between problem difficulty and target response length. Through follow-up preference optimization, we enable the slow-thinking model to acquire adaptive reasoning capabilities, strategically allocating more computational resources to challenging problems while maintaining efficient processing of simpler tasks. The proposed DAST method essentially establishes a learnable mapping between problem difficulty levels and corresponding target response length, achieving intelligent computation allocation during the inference stage without compromising reasoning quality.
@@ 2.1 Problem Difficulty Evaluation
Some previous studies have primarily relied on proprietary models such as ChatGPT to assess the difficulty or complexity of data samples [7, 6]. Such methods are limited by the capabilities of LLMs used for evaluation. In reasoning scenarios such as mathematics, a more common solution is to use sampling accuracy to measure the difficulty of the problem [13, 12]. However, this approach has two limitations. First, it requires sampling a larger number of answers to ensure effectiveness. Second, for some more difficult questions, there are fewer valid answers, resulting in inadequate discriminative power. The token length budget metric we propose in this paper can effectively circumvent these two limitations.
2.2 OverThinking Solution
Currently, two studies have addressed the phenomenon of overthinking in slow-thinking models and proposed corresponding solutions. [1] introduced a method based on length contrastive optimization approach, which substantially reduces the reasoning length. [14] developed a redundant token compression method to address this issue. Nevertheless, both approaches neglect problem difficulty differentiation, leading to substantial performance degradation on complex problems. In this paper, we propose a difficulty adaptive slow thinking approach to achieve a trade-off between reasoning length and performance.
@@ 3 Methodology
In this section, we introduce our proposed DAST method in detail. Our key insight lies in enhancing existing reasoning models through length-preference training, enabling adaptive response generation with length that corresponds to problem complexity. The main challenge lies in establishing a principled relationship between response length and problem difficulty. To this end, we propose a novel reasoning Token Length Budget (TLB) metric that dynamically scales with problem complexity: simpler questions receive smaller length allocations while complex ones are allocated extended budgets. This metric not only serves as a length reference for response generation but also could be used to quantify problem difficulty.
The technical implementation of DAST involves three crucial steps: First, calibrating the initial rule-based reward scores of each response with thinking process via comparing its actual token length with the TLB of the corresponding problem. Second, constructing a pairwise length-preference training dataset based on the calibrated reward scores. Finally, employing SimPO [9] to fine-tune the original reasoning model, endowing it with adaptive reasoning capabilities. The overall framework of our proposed method is depicted in Figure 2.
Figure 2: Overview of our proposed DAST method.
@@ 3.1 Token Length Budget Definition
denotes the sampling accuracy. Here, $c$ is the number of correct responses sampled from the current question, $N$ is the total number of sampled responses. $L_{\overline{r}}$ represents the average token length of correct responses, and $L_{\mathrm{max}}$ is the maximum generation length.
The higher the sampling accuracy, the closer $L_{budget}$ is to the average length of correct responses ($L_{\overline{r}}$ ), and the lower the sampling accuracy, the closer the $L_{budget}$ is to the maximum generated length. When the sampling accuracy is 0, it means that the current problem is extremely difficult, under these circumstances, the model should be encouraged to think as much as possible and generate longer COT. At this time, TLB is equal to the maximum generation length of the model. As shown in Figure 3, the average TLB exhibits strong positive correlation with problem difficulty level on the MATH training dataset, demonstrating its potential as an effective measure for quantifying problem complexity.
Figure 3: Average token budget distribution across difficulty levels ($L_{\mathrm{max}}=4096$). Results are computed using responses sampled from QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B-preview on the MATH training set. The higher TLB for DeepSeek-R1-Distill-Qwen-32B-preview stems from its structured output format containing both reasoning chains and final answers.
@@ 3.2 Reward Score Calibration
In reasoning scenarios such as mathematics and coding, o1-like slow thinking models typically employ rule-based rewards as feedback signals for training [2, 12]. In this work, traditional rule-based rewards are calibrated by incorporating the deviation between actual response length and the TLB metric. This calibration allows the reward score to jointly capture both difficulty-aware information and length characteristics, enabling difficulty-adaptive training.
@@
The calibrated reward score for response $i$ is defined as Equation 2. From Figure 4, we can derive the following insights:
Figure 4: Calibrated reward function.
For a correct answer, if its length exceeds TLB, it will result in a decay of the reward. The simpler the question, the smaller the TLB. If the current generated length significantly surpasses TLB, the reward decay will be more severe. Conversely, if the generated length is below TLB, the reward will be amplified to some extent, thereby encouraging the model to generate shorter answers within corresponding TLB.
@@
3.3 Length Preference Data Construction
For each question $x$, $N$ candidate responses are sampled with their TLB $L_{\mathrm{budget}}^{(x)}$ computed via Equation 1. Corresponding reward scores are then derived using Equation 2. Subsequently, all responses are ranked by the reward scores to constructed contrastive pairs $(x,y_{w},y_{l})$, where $y_{w}$ strictly dominates $y_{l}$ in reward by a minimum margin $\delta$. The maximum number of contrastive pairs per question is limited to a threshold $K$.
The resultant dataset $D_{pre}$ comprises two categories of contrastive pairs: (1) pairs where both samples generate correct answers but the positive instance demonstrates significantly shorter length; (2) pairs where both responses are incorrect yet the positive sample exhibits substantially longer answer length compared to its negative counterpart.
3.4 Length Preference Training
The constructed dataset $\mathcal{D}_{\mathrm{pre}}$ enables alignment of reasoning LLMs through Simple Preference Optimization (SimPO) [9]. We chose SimPO due to its characteristic of being more sensitive in controlling answer length. The optimization objective is formulated as:
$\mathcal{L}_{\mathrm{SimPO}}(\pi_{\theta})=-E_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$ (3)
@@ 4.1 Experimental Setup
Comparative experiments are conducted on two representative slow-thinking reasoning models:DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-32B. For both of them, we sample 20 answers for each question in the PRM12K MATH dataset [5] training set with sampling temperature $\tau=0.6$. The derived TLB metrics and calibrated rewards construct the SimPO training set $\mathcal{D}_{\mathrm{pre}}$ with reward margin $\delta=0.3$. Each question retains at most 4 contrastive pairs, yielding 18,231 training instances for SimPO training.
@@
Evaluation is performed on:
  • MATH-500 [5]: 500 competition mathematics problems from the MATH[3] test set
  • AIME-2024 [8]: 30 problems from the American Invitational Mathematics Examination, which tests mathematical problem solving across multiple areas
with maximum sequence length constrained to 32,768 tokens.
Baselines
Two methods from [1] addressing overthinking are reproduced:
  • SFTShortest: Supervised fine-tuning using only the shortest correct sampled response of each problem as training data
  • SimPOShortest: SimPO with contrastive instance pairs, which take the shortest correct sampled response of each problem as positive instance and the longest as negative instance
Models MATH500 AIME2024
Accuracy Tokens Accuracy Tokens
DeepSeek-R1-Distill-Qwen-7B 92.8 3506 60.0 7448
DeepSeek-R1-Distill-Qwen-7B-SFTShortest 91.8 2954 50.0 10757
DeepSeek-R1-Distill-Qwen-7B-SimPOShortest 87.8 970 33.3 2737
DeepSeek-R1-Distill-Qwen-7B-SimPODAST 92.6 2802 53.3 6337
DeepSeek-R1-Distill-Qwen-32B 94.4 3781 73.0 10955
DeepSeek-R1-Distill-Qwen-32B-SFTShortest 92.6 2401 60.0 9009
DeepSeek-R1-Distill-Qwen-32B-SimPOShortest 89.4 1181 36.7 2690
DeepSeek-R1-Distill-Qwen-32B-SimPODAST 96.4 2013 66.7 5341
Table 1: Main results
4.2 Main Results
@@
Method Level 1 Level 2 Level 3 Level 4 Level 5
SimPOShortest 69.3% 61.0% 66.1% 70.1% 70.8%
SimPODAST 53.7% 40.1% 42.0% 41.5% 36.6%
Table 2: Comparison of the relative reduction ratios of average token lengths between SimPOShortest and SimPODAST across different difficulty levels in MATH500
We also compared the accuracy and average token length based on DeepSeek-R1-Distill-Qwen-32B on MATH500 according to difficulty level. As shown in Figures 5(a) and 5(b), SimPODAST achieved the best accuracy on Level 4 and Level 5 with significant margin against other methods, which demonstrates that it maintains its reasoning ability under complex problems. Although SimPOShortest shows the most significant reduction in response length, its reasoning capability notably declines when addressing complex problems.
Statistical analysis of output length reduction ratios in Table 2 reveals that SimPOShortest exhibits limited differentiation in length reduction across different difficulty levels. In contrast, the DAST method shows discernible adaptive capabilities, achieving approximately 53.7% length reduction at Level 1 compared to the original model, while this reduction decreases to approximately 36.6% at the most challenging Level 5. This progressive performance variation validating its difficulty-adaptive nature.

### v2

# DAST: Difficulty-Adaptive Slow Thinking for Large Reasoning Models

@@ Abstract
Recent advancements in slow thinking reasoning models have shown exceptional performance in complex reasoning tasks. However, these models often exhibit overthinking (generating redundant reasoning steps for simple problems), leading to excessive computational resource usage. While current mitigation strategies uniformly reduce reasoning tokens, they risk degrading performance on challenging tasks that require extended reasoning. This paper introduces Difficulty-Adaptive Slow Thinking (DAST), a novel framework that enables models to autonomously adjust the length of Chain-of-Thought (CoT) based on problem difficulty. We first propose a Token Length Budget (TLB) metric to quantify difficulty, then leverage budget-aware reward shaping and budget preference optimization to implement DAST. DAST penalizes overlong responses for simple tasks while incentivizing sufficient reasoning for complex problems. Experiments on diverse datasets and model scales demonstrate that DAST effectively mitigates overthinking (reducing token usage by over 30% on average) while preserving reasoning accuracy on complex problems. Our codes and models are available at https://github.com/AnonymousUser0520/AnonymousRepo01.
DAST: Difficulty-Adaptive Slow Thinking for Large Reasoning Models
Yi Shen1,2, Jian Zhang1,2, Jieyun Huang1,2, 🖂Shuming Shi1,2, Wenjing Zhang1,2, Jiangze Yan1,2, Ning Wang1,2, Kai Wang1,2, Zhaoxiang Liu1,2, 🖂Shiguo Lian1,2 1Unicom Data Intelligence, China Unicom 2Data Science & Artificial Intelligence Research Institute, China Unicom 🖂 Corresponding Authors {sheny73, zhangj2791,liansg}@chinaunicom.cn, ssm01@hotmail.com
@@ 1 Introduction
Recently, significant advancements have been made in slow-thinking reasoning models, exemplified by OpenAI’s o1 OpenAI (2024) and DeepSeek-R1 Guo et al. (2025). These slow-thinking reasoning models, which simulate human deep-thinking mechanisms through self-reflection, error correction, and exploration of multiple solution strategies, have demonstrated remarkable potential in complex reasoning tasks such as mathematical problem-solving MAA (2024) and programming challenges Jain et al. (2024).
Figure 1: An Example to illustrate the overthinking phenomenon.
@@
However, empirical studies Chen et al. (2024); Sui et al. (2025); Liu et al. (2025) have shown that these reasoning models suffer from overthinking phenomenon. In other words, these models tend to generate redundant solutions and unnecessarily complex reasoning steps when addressing simple problems, leading to inefficient computational resource utilization. For instance, as demonstrated in Figure 1, traditional LLM (DeepSeek V3) can solve basic mathematic problems such as “3x + 7=22, x=?” with just 58 tokens, while reasoning model such as DeepSeek-R1 may consume over 1000 tokens for the same problem. This overthinking phenomenon not only significantly reduces reasoning efficiency but also causes information overload for the users.
@@
Current approaches Sui et al. (2025); Xia et al. (2025); Chen et al. (2024) to mitigate the overthinking problem typically employ a one-size-fits-all strategy, uniformly reducing reasoning steps or token counts across all problems. Although these approaches significantly reduce the output length of slow-thinking models, they carry the risk of performance degradation, particularly when addressing challenging problems. Prior studies Zeng et al. (2024); Muennighoff et al. (2025) have demonstrated that adequate reasoning length is critical for slow-thinking models to effectively solve complex tasks. It is therefore essential to devise approaches that mitigate overthinking phenomena while maximally preserving reasoning capabilities.
@@
This raises a fundamental question: Can slow thinking models autonomously adjust reasoning depth based on problem difficulty, thereby generating concise responses for simple questions while maintaining sufficiently extended CoT reasoning for complex ones? We propose a Difficulty-Adaptive Slow-Thinking (DAST) framework to tackle this challenge.
@@
Our key idea is straightforward: Given that tasks of varying difficulty levels inherently demand different reasoning depths, we propose to establish a mapping relationship between problem complexity and target response length. By comparing the length of the current response with the target response length, we can determine whether to apply additional rewards or penalties to the current answer. Building upon this, we construct a training objective to achieve adaptive reasoning. Specifically, we first introduce a difficulty quantification metric termed “Token Length Budget” (TLB), which integrates both the accuracy of sampled responses and their length distributions. This metric effectively combines problem difficulty characteristics with token length information. For multiple generated responses sampled, our method applies budget-aware reward shaping: Responses exceeding the TLB of simple questions receive penalty signals, while those approaching the TLB for complex problems receive positive incentives. This mechanism allows us to construct pair-wise budget preference training data that inherently encodes the relationship between problem difficulty and target response length. Through follow-up preference optimization, we enable the slow-thinking model to acquire adaptive reasoning capabilities, strategically allocating more computational resources to challenging problems while maintaining efficient processing of simpler tasks. The proposed DAST method essentially establishes a learnable mapping between problem difficulty levels and corresponding target response length, achieving intelligent computation allocation during the inference stage without compromising reasoning quality.
@@ 2.1 Problem Difficulty Evaluation
Some previous studies mainly used proprietary models like ChatGPT to assess the difficulty or complexity of data samples Lu et al. (2023); Liu et al. (2023). Such methods are limited by the ability of the LLMs they employed for evaluation. In reasoning scenarios such as mathematics, a more common solution is to use sampling accuracy to measure the difficulty of the problem Tong et al. (2025); Team et al. (2025). However, this approach has two shortcomings. First, it requires sampling more answers to ensure effectiveness. Second, for some extremely difficult questions, there are fewer valid answers, resulting in insufficient discrimination. The token length budget metric we propose in this paper can effectively circumvent these two shortcomings.
2.2 Overthinking
Mitigating “overthinking” in Large Reasoning Models (LRMs) to enhance reasoning efficiency has garnered increasing research attention. Existing approaches can be broadly categorized into three main types (Sui et al., 2025; Liu et al., 2025): prompt-based methods, output-based methods, and post-training methods.
Prompt-based Methods. These methods focus on modifying input prompts to achieve concise reasoning. Common techniques include imposing explicit token limits (Nayab et al., 2024) or instructing the model to generate more succinct reasoning chains (Xu et al., 2025; Renze and Guven, 2024).
Output-based Methods. These approaches intervene during the inference stage to control the reasoning process. One line of work compresses intermediate reasoning steps into latent representations rather than explicit text, thereby enhancing brevity at the cost of interpretability (Hao et al., 2024; Shen et al., 2025). Another direction involves dynamic decoding, wherein lightweight models or heuristics evaluate each reasoning step during generation, deciding whether to retain, modify, or discard it (Sun et al., 2024; Yang et al., 2025; Zhang et al., 2025).
Post-training Methods. These methods typically involve supervised fine-tuning (SFT) on variable-length Chain-of-Thought (CoT) data (Chen et al., 2024; Ma et al., 2025; Kang et al., 2025; Xia et al., 2025; Munkhbat et al., 2025) or the incorporation of length-based rewards within reinforcement learning frameworks (Team et al., 2025; Luo et al., 2025; Arora and Zanette, 2025; Yeo et al., 2025). The objective is to guide the model towards generating reasoning paths that are both concise and accurate.
While all these methods have demonstrated promising results in efficient reasoning, they exhibit certain limitations. Most existing approaches indiscriminately compress the Chain-of-Thought (CoT) across all problems, leading to degraded reasoning performance, particularly for complex problems. Furthermore, the majority of these methods have primarily been evaluated on LRMs with fewer than 7B parameters, with limited exploration of their efficacy on models with larger parameter scales (e.g., 32B). Our work aims to achieve difficulty-adaptive efficient inference by defining a token length budget for each problem, reflecting its perceived difficulty for the current model. Although some prior studies Aggarwal and Welleck (2025); Muennighoff et al. (2025) have employed token budgets to control inference length, these approaches typically rely on pre-defined, discrete token budgets, often manually set. The most closely related work is that of Han et al. (2024), which also attempts to allocate varying token budgets based on problem complexity. However, their method necessitates an iterative search process, experimenting with token limits within prompts to determine the final budget, which can be cumbersome. Additionally, its effectiveness has not been demonstrated on “slow thinking” models.
Figure 2: Overview of our proposed DAST method.
@@ 3 Methodology
In this section, we introduce our proposed DAST method in detail. Our key insight lies in enhancing existing reasoning models through budget-preference training, enabling adaptive response generation with length that corresponds to problem complexity. The main challenge lies in establishing a principled relationship between response length and problem difficulty. To this end, we propose a novel reasoning Token Length Budget (TLB) metric that dynamically scales with problem complexity: simpler questions receive smaller length allocations while complex ones are allocated extended budgets. This metric not only serves as a length reference for response generation but also could be used to quantify problem difficulty.
The technical implementation of DAST involves three crucial steps: First, calibrating the initial rule-based reward scores of each response with thinking process via comparing its actual token length with the TLB of the corresponding problem. Second, constructing a pairwise budget-preference training dataset based on the calibrated reward scores. Finally, employing SimPO Meng et al. (2025) to fine-tune the original reasoning model, endowing it with adaptive reasoning capabilities. The overall framework of DAST is depicted in Figure 2.
@@ 3.1 Token Length Budget Definition
denotes the sampling accuracy. Here, $c$ is the number of correct responses sampled from the current question with the backbone LRM, $N$ is the total number of sampled responses. $L_{\overline{r}}$ represents the average token length of correct responses, and $L_{\mathrm{max}}$ is the maximum generation length.
The higher the sampling accuracy, the closer $L_{budget}$ is to the average length of correct responses ($L_{\overline{r}}$), while lower accuracy brings $L_{budget}$ closer to the maximum generated length. A sampling accuracy of 0 indicates extreme difficulty, in which case the model should be encouraged to think deeply and generate longer CoT. At this point, TLB equals the model’s maximum generation length. As shown in Figure 3, the average TLB exhibits strong positive correlation with problem difficulty level on the MATH training dataset, demonstrating its potential as an effective measure for quantifying problem complexity.
@@ 3.2 Reward Score Calibration
In reasoning scenarios such as mathematics and coding, o1-like slow thinking models typically employ rule-based rewards as feedback signals for training Guo et al. (2025); Team et al. (2025). In this work, traditional rule-based rewards are calibrated by incorporating the deviation between actual response length and the TLB metric. This calibration allows the reward score to jointly capture both difficulty-aware information and length characteristics, enabling difficulty-adaptive training.
@@
The calibrated reward score for the response $i$ is defined as Equation 2. From Figure 4, we can derive the following insights:
For a correct answer, if its length exceeds TLB, it will result in a reward decay. The simpler the question, the smaller the TLB. If the generated length significantly surpasses TLB, the reward will decay severely. Conversely, if it falls below TLB, the reward will be amplified, encouraging the model to generate shorter answers within the corresponding TLB.
@@
Figure 3: Average token budget distribution across difficulty levels ($L_{\mathrm{max}}=4096$). Results are computed using responses sampled from QwQ-32B-preview and DeepSeek-R1-Distill-Qwen-32B (DS-32B) on the MATH training set. The higher TLB for DS-32B stems from its structured output format containing both reasoning chains and final answers.
3.3 Budget Preference Data Construction
For each input question $x$, $N$ candidate responses are sampled with corresponding TLB $L_{\mathrm{budget}}^{(x)}$ computed as formalized in Equation 1. The corresponding reward scores are then derived using Equation 2. These responses are subsequently ranked based on their reward scores to constructed contrastive pairs $(x,y_{w},y_{l})$ for subsequent preference optimization, where $y_{w}$ and $y_{l}$ denote the winning and losing responses respectively.
We categorize contrastive pairs into two distinct classes  [^1] There is actually a third class: Correct-InCorrect Pair (CICP), but our experiments show that CICP does not improve the performance. We will discuss the effects of CICP in Section 4.2.3 : (1) Dual-Correct Pair (DCP): Both responses yield correct answers, but the preferred instance $y_{w}$ demonstrates significantly higher output conciseness ($|y_{w}|\ll|y_{l}|$). DCP is designed to encourage the model to generate responses that are both correct and as concise as possible within the token length budget. (2) Dual-InCorrect Pair (DICP): Both responses produce incorrect answers, yet $y_{w}$ exhibits substantially longer reasoning chains ($|y_{w}|\gg|y_{l}|$). DICP is designed to stimulate more extensive reasoning attempts when the model has not yet produced a correct answer and remains within the corresponding TLB.
For each question, we first select the DCP and DICP pairs with maximal reward margin $\Delta R=R(y_{w})-R(y_{l})$, then apply a two-stage filtering process: 1. We establish a truncation threshold $\delta\in(0,1)$ to eliminate the bottom $\delta|D|$ pairs with minimal $\Delta R$, where $|D|$ denotes the candidate set size. 2. To maintain data quality and training efficiency, we retain at most two highest-margin pairs (one DCP and one DICP) per question.
This selection mechanism ensures statistical significance in reward differences while preserving informative contrastive signals, ultimately enhancing the stability of preference optimization.
Figure 4: Calibrated reward function with TLB.
3.4 Budget Preference Training
The constructed dataset $\mathcal{D}_{\mathrm{pre}}$ enables alignment of reasoning LLMs through Simple Preference Optimization (SimPO) Meng et al. (2025). We chose SimPO due to its characteristic of being more sensitive in controlling answer length. The optimization objective is formulated as:
$\mathcal{L}_{\mathrm{SimPO}}(\pi_{\theta})=-E_{(x,y_{w},y_{l})\sim\mathcal{D}}\Bigl{[}\log\sigma\Bigl{(}\\ \frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\Bigr{)}\Bigr{]},$ (3)
@@ 4.1 Experimental Setup
Backbone Reasoning Models
We conduct comparative experiments on two Large Reasoning Models (LRMs): DeepSeek-R1-DistillQwen-7B (DS-7B) and DeepSeek-R1-Distill-Qwen-32B (DS-32B) Guo et al. (2025). Although both models exhibit strong reasoning abilities, their substantial redundant thought processes highlight the necessity of response compression.
@@
We evaluate model performance on three established reasoning benchmarks:
  • MATH-500 Lightman et al. (2023): 500 high school competition-level mathematical problems stratified into 5 difficulty levels, ranging from Level 1 (easiest) to Level 5 (hardest);
  • AIME 2024 MAA (2024): 30 curated problems from the American Invitational Mathematics Examination testing complex problem-solving;
  • GPQA Rein et al. (2024): 198 PhD-level science questions across physics, chemistry, and biology.
Baseline Methods
We compare our method against the following representative approaches designed for efficient reasoning:
  • Concise Thoughts (CCoT) Renze and Guven (2024): It encourages the model to generate concise reasoning process via simply append “Be concise” to the prompt.
  • Chain of Draft (CoD) Xu et al. (2025): This is another prompt-based method which instructs the model to generate concise draft intermediate steps during reasoning.  [^2] Please refer to the Appendix 8 for specific prompt templates for CCoT and CoD.
  • SFTShortest Munkhbat et al. (2025); Chen et al. (2024): This method selects the shortest correct response from the backbone model’s sampled answers as the ground truth, and then performs supervised fine-tuning (SFT) on the backbone model.
  • SimPOShortest Chen et al. (2024): SimPO with contrastive instance pairs generated by the backbone reasoning model, which takes the shortest correct sampled response of each problem as positive instance and the longest correct counterpart as negative instance.
  • SimPOCosine: We keep the DAST settings entirely unchanged, only replacing the ranking criterion for contrastive pairs from the reward function defined in Section 3.2 to the cosine reward function introduced in Yeo et al. (2025). We aim to verify the effectiveness of our proposed reward function through comparative analysis with SimPOCosine.
  • SimPOLenPenalty: We employ the length penalty reward function defined in Team et al. (2025) to evaluate the sampled responses for each question, select the highest and lowest ranked instances to construct contrastive pairs, and thereby develop another version of SimPO baseline.
Evaluation Metrics
We adopt following metrics to comprehensively assess both reasoning accuracy and compression effectiveness: ACC denotes the accuracy of the final answer. LEN refers to the average response length, measured in tokens. C-LEN represents the average number of tokens in all correct responses. CR is the compression ratio, which is defined as token length reduction ratio (vs. original model). C-CR is the C-LEN reduction ratio against original model.
Training Details
For both backbone reasoning models, we generate 20 candidate responses for each question in the MATH Hendrycks et al. (2021) training set with maximum sequence length constrained to 4,096 tokens to compute its TLB. Following reward score calibration via Equation 2, we construct the preference training set for SimPO optimization. The truncation threshold $\delta$ is set to 0.15 and 0.18 for DS-7B and DS-32B, yielding final training sets of 10295 and 9813 contrastive pairs for DS-7B and DS-32B, respectively. All models are trained for 1 epoch using AdamW optimizer with learning rate lr = 5e-6. All our experiments were run on a NVIDIA GPU machine with 8 $\times$ H100. Please refer to Appendix A for complete training configuration statistics.
Decoding Configuration
In our evaluation setup, all models were constrained to a maximum generation length of 32,768 tokens to align with DeepSeek’ technical report Guo et al. (2025). Following Chen et al. (2024); Yeo et al. (2025), we employ greedy decoding for all the models. Results were computed using OpenR1 evaluation scripts [^3] https://github.com/huggingface/open-r1.
MODEL METHOD MATH-500 AIME 2024 GPQA
@@
DS-7B Origin 93.2 4039 3506 - - 60.0 10603 7448 - - 47.98 8021 7242 - -
CCoT 92.2 3388 2887 16.1% 17.7% 40.0 10976 6497 -3.5% 12.8% 47.98 7612 6914 5.1% 4.5%
CoD 75.8 1596 1234 60.5% 64.8% 43.3 9399 6630 11.4% 11.0% 49.49 7178 6932 10.5% 4.3%
SimPOshortest 89.8 1891 1557 53.2% 55.6% 53.3 8291 3847 21.8% 48.4% 50.51 6068 5401 24.4% 25.4%
SFTshortest 91.8 2987 2408 26.0% 31.3% 50.0 12989 6227 -22.5% 16.4% 51.01 7760 6578 3.3% 9.2%
SimPOcosine 93.2 3897 3223 3.5% 8.1% 63.3 12572 6783 -18.6% 8.9% 50.00 8230 6912 -2.6% 4.6%
SimPOLenPenalty 89.4 1922 1612 52.4% 54.0% 63.3 7419 4478 30.0% 39.9% 51.01 5860 4847 26.9% 33.1%
DAST(ours) 93.6 3309 2709 18.1% 22.8% 70.0 10804 7924 -1.9% 6.4% 51.51 7684 6480 4.2% 10.5%
DS-32B Origin 94.4 3782 3384 - - 73.3 10955 9124 - - 65.15 6410 5923 - -
CCoT 93.2 2044 1733 46.0% 48.8% 56.7 8436 5678 23.0% 37.8% 63.13 5820 5143 9.2% 13.2%
CoD 93.6 1941 1628 48.7% 51.9% 43.3 7288 5065 33.5% 44.5% 62.12 5107 4831 20.3% 18.4%
SimPOshortest 89.0 1107 998 70.7% 70.5% 36.7 2580 855 76.4% 90.6% 63.13 2455 2286 61.7% 61.4%
SFTshortest 94.6 2402 2141 36.5% 36.7% 66.7 8204 6577 25.1% 27.9% 64.65 6044 5030 5.7% 15.1%
SimPOcosine 94.2 2325 1968 38.5%  41.8% 63.3 7379 5317 32.6% 41.7% 65.15 5835 4987 9.0% 15.8%
SimPOLenPenalty 90.6 1190 1066 68.5% 68.5% 43.3 2748 2047 74.9% 77.6% 62.12 2375 2111 62.9% 64.4%
DAST(ours) 95.8 2044 1744 46.0% 48.5% 76.7 7023 5409 35.9% 40.7% 65.15 5535 4514 13.7% 23.8%
Table 1: Evaluation results across the benchmarks.
4.2 Results and Analysis
4.2.1 Overall Results
The main results are presented in Table 1. We have the following findings: Prompt-based methods (CCoT, CoD) show unstable performance, often incurring accuracy losses, which are particularly pronounced on complex task AIME 2024. For example, ACC of CoD with DS-7B on AIME 2024 drops from 60.0% to 43.3%. Aggressive compression methods (SimPOShortest, SimPOLenPenalty) achieve the most significant token reduction on both DS-7B and DS-32B across all benchmarks. However, this substantial compression invariably sacrifices some accuracy. SimPOLenPenalty demonstrates a slightly better overall balance against SimPOShortest , potentially because its reward function introduces the average length of batch data as a comparison baseline, thereby better navigating the length-accuracy trade-off. SFTShortest proves to be a strong baseline, especially with the DS-32B model. However, notably on the AIME 2024, SFTShortest fails to reduce output length effectively, it is plausible that the straightforward SFT with shortest responses may have compromised the model’s instruction following ability, resulting in ineffective termination of responses when confronted with complex tasks such as AIME 2024.
DAST and SimPOCosine exhibit similar overall trends in balancing ACC and CR, the potential reason may be that neither method strictly prioritizes brevity but can encourage longer responses when beneficial. The superior performance of DAST over the standard cosine-based reward across benchmarks on both ACC and CR validates the effectiveness of our proposed budget-based reward function.
Despite being exclusively math-trained, DAST (DS-7B) achieves 51.51% (+3.53%) on GPQA with modest compression ratio (4.2%), demonstrating certain ability of domain generalization. On the challenging AIME 2024 benchmark, DAST (DS-7B) does not reduce the average response length (CR -1.9%). This, combined with a substantial accuracy improvement from 60.0% (Origin) to 70.0% suggests that DAST does not indiscriminately shorten reasoning paths but can adaptively allocate more reasoning steps for complex problems.
Overall, the results in Table 1 affirm that DAST effectively navigates the intricate trade-off between conciseness and reasoning performance. It generally preserves or improves the reasoning capabilities of the backbone models while achieving remarkable CoT reductions, outperforming the baselines in this combined objective. This is particularly evident with the more capable DS-32B model, where DAST achieves strong compression alongside ACC improvements across all the benchmarks.